# Maximal discs of Weil-Petersson class in $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$

### v1

@@ 5. Variation of anti-holomorphic energy
In this section, we prove that the anti-holomorphic energy is a potential function of the symplectic form $\omega_{C}$ restricted to the subspaces $T_{0}(1)^{\pm}$ of $T^{*}T_{0}(1)$. Recall that $T_{0}(1)^{+}$, $T_{0}(1)^{-}$ denote the inverse image of $\mathrm{Mess}$ of $T_{0}(1)\times\{0\}$, $\{0\}\times T_{0}(1)$ respectively.
First, we consider a variation for the holomorphic energy density and anti-holomorphic energy density.
@@
For a family of anti-holomorphic energy densities $e^{\psi^{\epsilon}\circ F^{\epsilon}}|F^{\epsilon}_{\bar{z}}|^{2}$ on $\Sigma^{\epsilon}$ satisfying the commutative diagram (4.1),
@@
From $h^{\epsilon}\circ F=F^{\epsilon}\circ f^{\epsilon}$, we have
@@
Combining (5.2) and (5.3),
@@
$=$ $\ e^{\psi\circ F}F_{z}\overline{F}_{z}\Big{(}\dot{h}_{\bar{z}}\circ F\overline{F}_{\bar{z}}F_{z}^{-1}-\dot{f}_{\bar{z}}\Big{)}+e^{\psi\circ F}F_{\bar{z}}\overline{F}_{\bar{z}}\Big{(}\dot{\bar{h}}_{z}\circ FF_{z}\overline{F}_{\bar{z}}^{-1}-\dot{\bar{f}}_{z}\Big{)}$
@@
For the third equality above, we used the equality (4.13) for a variation of the hyperbolic metrics as before. This completes the proof. ∎
In the same way as the above, we can prove
@@
For a family of holomorphic energy densities $e^{\phi^{\epsilon}}=e^{\psi^{\epsilon}\circ F^{\epsilon}}|F^{\epsilon}_{z}|^{2}$ on $\Sigma^{\epsilon}$ satisfying the commutative diagram (4.1),
@@
Recall that $T_{0}(1)^{\pm}$ is a symplectic space with the restriction of $\omega_{\mathrm{C}}$, which is the same as $\mp\mathrm{Mess}^{*}(\omega_{\mathrm{WP}})$ by Theorem 4.10. Using (2.9), along $T_{0}(1)^{\pm}$, we can derive
(5.5) $\mu_{z_{\pm}}=\frac{2\mu_{z}}{1+|\mu_{z}|^{2}},\qquad\mu_{F_{+}}=-(z^{-1})^{*}(\mu_{z})$
where $\mu_{z_{\pm}}$, $\mu_{z}$ denotes the Beltrami differential of the map $z_{\pm}$, $z$ from $(\mathbb{D},w)$.
@@
Under the condition $F_{+}^{*}(\nu_{+})=0$ or $F_{-}^{*}(\nu_{-})=0$,
(5.6) $\frac{1}{2}\big{(}L_{\nu}-iL_{i\nu}\big{)}\big{(}e^{\psi\circ F}|F_{\bar{z}}|^{2}\big{)}=e^{\phi}\bar{\mu}_{F}\nu.$
@@
We will prove only for the case $F_{-}^{*}(\nu_{-})=0$ since the proof for the other case is the same. By (5.1),
@@
Under the condition $F_{+}^{*}(\nu_{+})=0$ or $F_{-}^{*}(\nu_{-})=0$,
(5.9) $\frac{1}{2}\big{(}L_{\mu}+iL_{i\mu}\big{)}\big{(}e^{\phi}\nu\bar{\mu}_{F}\big{)}=e^{\phi}(1+|\mu_{F}|^{2})\nu\bar{\mu}.$
@@
We will prove only for the case $F_{-}^{*}(\nu_{-})=0$ since the proof for the other case is the same. By (4.10) and $L_{\mu}\nu=0$ (see (2.3) of [20]),
$L_{\mu}\big{(}e^{\phi}\nu\bar{\mu}_{F}\big{)}=L_{\mu}\big{(}\Phi\nu\big{)}$
@@
Hence,
$\frac{1}{2}\big{(}L_{\mu}+iL_{i\mu}\big{)}\big{(}e^{\phi}\nu\bar{\mu}_{F}\big{)}=e^{\phi}(1+|\mu_{F}|^{2})\nu\bar{\mu}.$
@@
(5.10) ¯$\partial\bar{\partial}\Big{(}2\int_{\Sigma}e^{\psi\circ F}|F_{\bar{z}}|^{2}\,d^{2}z\Big{)}=\mp\sqrt{-1}\,i^{*}_{\pm}\omega_{\mathrm{C}},$
where $i_{\pm}:T_{0}(1)^{\pm}\to T^{*}T_{0}(1)$ denotes the embedding map of $T_{0}(1)^{\pm}$ respectively.
@@
We prove only for the case of $T_{0}(1)^{+}$ since the proof of the other case is essentially same.
In order to consider variation of the anti-holomorphic energy of $F_{\pm}$, we have to deal with variation of the anti-holomorphic density $e^{\phi}|\mu_{F}|^{2}\,d^{2}z$ together. For the variation of $d^{2}z=-\sqrt{-1}/2dz\wedge d\bar{z}$ along a family of quasi-conformal map $f^{\epsilon}$, we have
$df^{\epsilon}\wedge d\bar{f^{\epsilon}}=\big{(}|f^{\epsilon}_{z}|^{2}-|f^{\epsilon}_{\bar{z}}|^{2}\big{)}\,dz\wedge d\bar{z}=|f^{\epsilon}_{z}|^{2}\big{(}1-\epsilon^{2}|\mu_{f}|^{2}\big{)}\,dz\wedge d\bar{z}$
where we used the equality $f^{\epsilon}_{\bar{z}}=\epsilon\mu_{f}f^{\epsilon}_{z}$ for the second equality. Hence, the factor $|f^{\epsilon}_{z}|^{2}$ with the variational term of $e^{\phi}|\mu_{F}|^{2}$ precisely gives the Lie derivative of this in Propositions 5.3 and 5.4. Then, we have
(5.11) $\begin{split}\partial_{\nu}\Big{(}2\int_{\Sigma}e^{\psi\circ F}|F_{\bar{z}}|^{2}\,d^{2}z\Big{)}=&2\int_{\Sigma}e^{\phi}\,\nu\bar{\mu}_{F}\,d^{2}z,\\ \bar{\partial}_{\mu}\partial_{\nu}\Big{(}2\int_{\Sigma}e^{\psi\circ F}|F_{\bar{z}}|^{2}\,d^{2}z\Big{)}=&2\int_{\Sigma}e^{\phi}\,(1+|\mu_{F}|^{2})\,\nu\bar{\mu}\,d^{2}z.\end{split}$
@@
(5.12) $\begin{split}\omega_{\mathrm{C}}=&-\frac{\sqrt{-1}}{2}\int_{\Sigma}(1-|\mu_{F}|^{2})\,e^{\phi}\,\sum_{k\in\mathbb{Z}}\Big{(}F^{*}_{+}(\nu_{+,k})\wedge F^{*}_{+}(\bar{\nu}_{+,k})\Big{)}\,d^{2}z\\ =&-\frac{\sqrt{-1}}{2}\int_{\Sigma}(1-|\mu_{F}|^{2})\,e^{\psi\circ F_{+}}|(F_{+})_{z}|^{2}\,\sum_{k\in\mathbb{Z}}\Big{(}F^{*}_{+}(\nu_{+,k})\wedge F^{*}_{+}(\bar{\nu}_{+,k})\Big{)}\,d^{2}z\\ =&-\frac{\sqrt{-1}}{2}\mathrm{Mess}^{*}_{+}\int_{\mathbb{D}}\,e^{\psi}\,\sum_{k\in\mathbb{Z}}\Big{(}\nu_{+,k}\wedge\bar{\nu}_{+,k}\Big{)}\,d^{2}z\\ =&-\frac{\sqrt{-1}}{2}\mathrm{Mess}^{*}_{+}\int_{\mathbb{D}}\,e^{\psi}\,\sum_{k,\ell\in\mathbb{Z}}\Big{(}\nu_{+,k}\wedge\bar{\nu}_{+,\ell}\Big{)}\,d^{2}z.\end{split}$
@@
(5.13) $\begin{split}&-\frac{\sqrt{-1}}{2}\mathrm{Mess}^{*}_{+}\int_{\mathbb{D}}\,e^{\psi}\,\Big{(}\nu_{+,k}\wedge\bar{\nu}_{+,\ell}+\nu_{+,\ell}\wedge\bar{\nu}_{+,k}\Big{)}\,d^{2}z\\ =&-\frac{\sqrt{-1}}{2}\int_{\Sigma}(1-|\mu_{F}|^{2})\,e^{\phi}\\ &\qquad\cdot\Big{(}2(1-|\mu_{F}|^{2})^{-1}(\nu_{k}-\bar{\nu}_{k}\mu_{F}^{2})\wedge 2(1-|\mu_{F}|^{2})^{-1}(\bar{\nu}_{\ell}-\nu_{\ell}\bar{\mu}_{F}^{2})\Big{)}\,d^{2}z\\ &-\frac{\sqrt{-1}}{2}\int_{\Sigma}(1-|\mu_{F}|^{2})\,e^{\phi}\\ &\qquad\cdot\Big{(}2(1-|\mu_{F}|^{2})^{-1}(\nu_{\ell}-\bar{\nu}_{\ell}\mu_{F}^{2})\wedge 2(1-|\mu_{F}|^{2})^{-1}(\bar{\nu}_{k}-\nu_{k}\bar{\mu}_{F}^{2})\Big{)}\,d^{2}z\\ =&-2{\sqrt{-1}}\int_{\Sigma}(1-|\mu_{F}|^{2})^{-1}\,e^{\phi}\\ &\qquad\qquad\cdot\Big{(}\nu_{k}\wedge\nu_{\ell}+\nu_{\ell}\wedge\bar{\nu}_{k}+\bar{\nu}_{k}\wedge\nu_{\ell}\,|\mu_{F}|^{4}+\bar{\nu}_{\ell}\wedge\nu_{k}\,|\mu_{F}|^{4}\Big{)}\,d^{2}z\\ =&-2\sqrt{-1}\int_{\Sigma}(1+|\mu_{F}|^{2})\,e^{\phi}\,\Big{(}\nu_{k}\wedge\bar{\nu}_{\ell}+\nu_{\ell}\wedge\bar{\nu}_{k}\Big{)}\,d^{2}z.\end{split}$
@@
(5.14) $-\sqrt{-1}\,i^{*}_{+}\omega_{C}=-2\int_{\Sigma}(1+|\mu_{F}|^{2})\,e^{\phi}\,\sum_{k,\ell\in\mathbb{Z}}\Big{(}\nu_{k}\wedge\bar{\nu}_{\ell}\Big{)}\,d^{2}z.$
@@
¯$\partial\bar{\partial}\Big{(}2\int_{\Sigma}e^{\phi}|\mu_{F}|^{2}\,d^{2}z\Big{)}=-\sqrt{-1}\,i^{*}_{+}\omega_{\mathrm{C}}\qquad\text{over}\quad T_{0}(1)^{+}.$
@@
By Theorem 4.10 and Theorem 5.5,
@@
¯$\partial\bar{\partial}\Big{(}2\int_{\Sigma}e^{\phi}|\mu_{F}|^{2}\,d^{2}z\Big{)}=\sqrt{-1}\,i^{*}_{\pm}\mathrm{Mess}^{*}_{\pm}\big{(}\omega_{\mathrm{WP}}\big{)}.$
@@
By Remark 4.12 and the proof of Theorem 5.5, the pullback 2-form $\mathrm{Mess}^{*}_{\pm}(\omega_{\mathrm{WP}})$ to $T_{0}(1)^{\pm}$ has the following expression
(5.15) $i^{*}_{\pm}\mathrm{Mess}^{*}_{\pm}(\omega_{\mathrm{WP}})=2{\sqrt{-1}}\int_{\Sigma}(1+|\mu_{F}|^{2})e^{\phi}\,\sum_{k,\ell\in\mathbb{Z}}\,\nu_{k}\wedge\bar{\nu}_{\ell}\ d^{2}z$

### v2

@@ 5. Variation of anti-holomorphic energy
In this section, we prove that the anti-holomorphic energy functional $E$ serves as a Kähler potential function for the canonical symplectic form $\omega_{C}$ when restricted to the submanifolds $T_{0}(1)^{\pm}$ of $T^{*}T_{0}(1)$. Although $T_{0}(1)^{\pm}$ are real symplectic submanifolds of the complex manifold $T^{*}T_{0}(1)$, they are endowed with a complex structure by identifying them with $T_{0}(1)$. This identification allows us to interpret $E$ as a Kähler potential with respect to the induced Kähler structure on $T_{0}(1)^{\pm}$.
Throughout the remainder of this section, we slightly abuse notation by writing $F=F_{\pm}$ whenever no confusion arises. We begin by considering variations of the holomorphic and anti-holomorphic energy densities.
@@
For a family of anti-holomorphic energy densities $e^{\psi^{\epsilon}\circ F^{\epsilon}}|F^{\epsilon}_{\bar{z}}|^{2}$ on $\mathbb{D}^{\epsilon}$ satisfying the commutative diagram (4.2),
@@
From $h^{\epsilon}\circ F=F^{\epsilon}\circ f^{\epsilon}$, we obtain
@@
By combining these expressions with the previous equation, we get
@@
Using the equality (4.13) for a variation of hyperbolic metrics, we conclude
$L_{{\nu_{f}}}\,\big{(}\,e^{\psi\circ F}|F_{\bar{z}}|^{2}\,\big{)}=$ $\ e^{\psi\circ F}F_{z}\overline{F}_{z}\Big{(}\dot{h}_{\bar{z}}\circ F\overline{F}_{\bar{z}}F_{z}^{-1}-\dot{f}_{\bar{z}}\Big{)}+e^{\psi\circ F}F_{\bar{z}}\overline{F}_{\bar{z}}\Big{(}\dot{\bar{h}}_{z}\circ FF_{z}\overline{F}_{\bar{z}}^{-1}-\dot{\bar{f}}_{z}\Big{)}$
@@
This completes the proof. ∎
In the same way as above, we can prove the following proposition.
@@
For a family of holomorphic energy densities $e^{\phi^{\epsilon}}=e^{\psi^{\epsilon}\circ F^{\epsilon}}|F^{\epsilon}_{z}|^{2}$ on $\mathbb{D}^{\epsilon}$ satisfying the commutative diagram (4.2), the following equality holds:
@@
For the following two propositions, we assume that either $F^{*}_{+}(\nu_{+})$ or $F^{*}_{-}(\nu_{-})$ vanishes, holding respectively over $T_{0}(1)^{-}$ and $T_{0}(1)^{+}$.
@@
Under the condition $F_{+}^{*}(\nu_{+})=0$ or $F_{-}^{*}(\nu_{-})=0$, we have
(5.5) $\frac{1}{2}\big{(}L_{\nu}-iL_{i\nu}\big{)}\big{(}e^{\psi\circ F}|F_{\bar{z}}|^{2}\big{)}=e^{\phi}\bar{\mu}_{F}\nu=\Phi\,\nu$
@@
We will prove the statement for the case $F_{-}^{*}(\nu_{-})=0$ as the proof for the other case follows similarly. From (5.1), we obtain
@@
Under the condition $F_{+}^{*}(\nu_{+})=0$ or $F_{-}^{*}(\nu_{-})=0$, we have
(5.8) $\frac{1}{2}\big{(}L_{\mu}+iL_{i\mu}\big{)}\big{(}e^{\phi}\bar{\mu}_{F}\,\nu\big{)}=e^{\phi}(1+|\mu_{F}|^{2})\nu\bar{\mu}.$
@@
We will prove the statement for the case $F_{-}^{*}(\nu_{-})=0$ as the proof for the other case follows similarly. From (4.10) and $L_{\mu}\nu=0$ (see (2.3) of [23]), we obtain
$L_{\mu}\big{(}e^{\phi}\bar{\mu}_{F}\,\nu\big{)}=L_{\mu}\big{(}\Phi\nu\big{)}$
@@
Hence, we obtain
$\frac{1}{2}\big{(}L_{\mu}+iL_{i\mu}\big{)}\big{(}e^{\phi}\bar{\mu}_{F}\,\nu\big{)}=e^{\phi}(1+|\mu_{F}|^{2})\nu\bar{\mu}.$
This completes the proof. ∎
Remark 5.5.
A point in the space $T_{0}(1)\times\{0\}$ or $\{0\}\times T_{0}(1)$ corresponds to a maximal conformal embedding $\sigma:\mathbb{D}_{z}\to\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$ such that one of the target discs of $F_{\pm}$ is fixed to be the origin, that is, $\mathbb{D}_{w}$. Equivalently, this means that one of $\mathbb{D}_{z_{\pm}}$ coincides with $\mathbb{D}_{w}$ in the commutative diagram (3.10). Hence, using (2.10), we can derive the following identities along $T_{0}(1)^{\pm}$:
(5.9) $\mu_{z_{\pm}}=\frac{2\mu_{z}}{1+|\mu_{z}|^{2}},\qquad\mu_{F_{\pm}}=-(z^{-1})^{*}(\mu_{z})$
where $\mu_{z_{\pm}}$, $\mu_{z}$ denotes the Beltrami differential of the maps $z_{\pm}$ and $z$ from $\mathbb{D}_{w}$.
Remark 5.6.
By construction, the submanifolds $T_{0}(1)^{\pm}\subset T^{*}T_{0}(1)$ consist of pairs $(\mu,\Phi)\in T^{*}T_{0}(1)$ where $\Phi$ is the Hopf differential of the harmonic map $F_{\pm}$ from the disc $\mathbb{D}_{z}$, determined $\mu$, to a fixed target disc $\mathbb{D}_{z_{\mp}}=\mathbb{D}_{w}$, as noted in Remark 5.5. This geometric setup implies that as $\mu$ varies in $T_{0}(1)$, the corresponding $\Phi$ depend only on $\mu$, thereby defining a section of the cotangent bundle $T^{*}T_{0}(1)$. The images of these sections are precisely the submanifolds $T_{0}(1)^{\pm}$.
However, such a section is not a holomorphic section, as the anti-holomorphic derivative of $\Phi$ does not vanish. In fact, one can easily verify that the image of a holomorphic section of $T^{*}T_{0}(1)$ is necessarily a Lagrangian submanifold with respect to $\omega_{\mathrm{C}}$. In contrast, $T_{0}(1)^{\pm}$ are real symplectic submanifolds of $T^{*}T_{0}(1)$, equipped with the restriction of $\omega_{\mathrm{C}}$.
Beyond the case discussed in [18] and referenced in the Introduction, a similar phenomenon occurs in the finite dimensional Teichmüller space setting, as explored in [20] and [17]. In those works, the corresponding submanifolds of the cotangent bundle arises as the differences between the Fuchsian projective structure and either the Schottky or Bers projective structures.
Finally, we show that the anti-holomorphic energy functional of the harmonic map $F_{\pm}$ is a Kähler potential function of the canonical symplectic form $\omega_{\mathrm{C}}$ over $T_{0}(1)^{\pm}\subset T^{*}T_{0}(1)$:
@@
(5.10) ¯$\partial\bar{\partial}\,\big{(}2E\big{)}=\partial\bar{\partial}\Big{(}2\int_{\mathbb{D}}e^{\psi\circ F}|F_{\bar{z}}|^{2}\,d^{2}z\Big{)}=\mp\sqrt{-1}\,i^{*}_{\pm}\omega_{\mathrm{C}},$
where $i_{\pm}:T_{0}(1)^{\pm}\to T^{*}T_{0}(1)$ denotes the embedding map of $T_{0}(1)^{\pm}$, respectively.
@@
We prove the case for $T_{0}(1)^{+}$ as the proof for the other case follows similarly.
To analyze the variation of the anti-holomorphic energy of $F_{\pm}$, we consider the variation of the anti-holomorphic density $e^{\phi}|\mu_{F}|^{2}\,d^{2}z$ along a family of quasi-conformal map $f^{\epsilon}$. This is given by
$-\frac{\sqrt{-1}}{2}\,{e^{\phi^{\epsilon}}\circ f^{\epsilon}\,|\mu_{F^{\epsilon}}|^{2}\circ f^{\epsilon}\,df^{\epsilon}\wedge d\bar{f^{\epsilon}}}$
$=$ $-\frac{\sqrt{-1}}{2}\,e^{\phi^{\epsilon}}\circ f^{\epsilon}\,|\mu_{F^{\epsilon}}|^{2}\circ f^{\epsilon}\,\big{(}|f^{\epsilon}_{z}|^{2}-|f^{\epsilon}_{\bar{z}}|^{2}\big{)}\,dz\wedge d\bar{z}$
$=$ $-\frac{\sqrt{-1}}{2}\,e^{\phi^{\epsilon}}\circ f^{\epsilon}\,|\mu_{F^{\epsilon}}|^{2}\circ f^{\epsilon}\,|f^{\epsilon}_{z}|^{2}\big{(}1-\mathrm{O}(\epsilon^{2})|\mu_{f}|^{2}\big{)}\,dz\wedge d\bar{z}$
where we used the equality $f^{\epsilon}_{\bar{z}}=\mu_{f}^{\epsilon}f^{\epsilon}_{z}$ for the last equality. Thus, the factor $|f^{\epsilon}_{z}|^{2}$ together with the variational term of $e^{\phi}|\mu_{F}|^{2}$ precisely corresponds to the Lie derivative described in Propositions 5.3 and 5.4. Consequently, we obtain
[…]
@@
(5.12) $\begin{split}\omega_{\mathrm{C}}=&-\frac{\sqrt{-1}}{2}\int_{\mathbb{D}}(1-|\mu_{F}|^{2})\,e^{\phi}\,\sum_{k\in\mathbb{Z}}\Big{(}F^{*}_{+}(\nu_{+,k})\wedge F^{*}_{+}(\bar{\nu}_{+,k})\Big{)}\,d^{2}z\\ =&-\frac{\sqrt{-1}}{2}\int_{\mathbb{D}}(1-|\mu_{F}|^{2})\,e^{\psi\circ F_{+}}|(F_{+})_{z}|^{2}\,\sum_{k\in\mathbb{Z}}\Big{(}F^{*}_{+}(\nu_{+,k})\wedge F^{*}_{+}(\bar{\nu}_{+,k})\Big{)}\,d^{2}z\\ =&-\frac{\sqrt{-1}}{2}\mathrm{Mess}^{*}_{+}\int_{\mathbb{D}}\,e^{\psi}\,\sum_{k\in\mathbb{Z}}\Big{(}\nu_{+,k}\wedge\bar{\nu}_{+,k}\Big{)}\,d^{2}z\\ =&-\frac{\sqrt{-1}}{2}\mathrm{Mess}^{*}_{+}\int_{\mathbb{D}}\,e^{\psi}\,\sum_{k,\ell\in\mathbb{Z}}\Big{(}\nu_{+,k}\wedge\bar{\nu}_{+,\ell}\Big{)}\,d^{2}z.\end{split}$
@@
(5.13) $\begin{split}&-\frac{\sqrt{-1}}{2}\mathrm{Mess}^{*}_{+}\int_{\mathbb{D}}\,e^{\psi}\,\Big{(}\nu_{+,k}\wedge\bar{\nu}_{+,\ell}+\nu_{+,\ell}\wedge\bar{\nu}_{+,k}\Big{)}\,d^{2}z\\ =&-\frac{\sqrt{-1}}{2}\int_{\mathbb{D}}(1-|\mu_{F}|^{2})\,e^{\phi}\\ &\qquad\cdot\Big{(}2(1-|\mu_{F}|^{2})^{-1}(\nu_{k}-\bar{\nu}_{k}\mu_{F}^{2})\wedge 2(1-|\mu_{F}|^{2})^{-1}(\bar{\nu}_{\ell}-\nu_{\ell}\bar{\mu}_{F}^{2})\Big{)}\,d^{2}z\\ &-\frac{\sqrt{-1}}{2}\int_{\mathbb{D}}(1-|\mu_{F}|^{2})\,e^{\phi}\\ &\qquad\cdot\Big{(}2(1-|\mu_{F}|^{2})^{-1}(\nu_{\ell}-\bar{\nu}_{\ell}\mu_{F}^{2})\wedge 2(1-|\mu_{F}|^{2})^{-1}(\bar{\nu}_{k}-\nu_{k}\bar{\mu}_{F}^{2})\Big{)}\,d^{2}z\\ =&-2{\sqrt{-1}}\int_{\mathbb{D}}(1-|\mu_{F}|^{2})^{-1}\,e^{\phi}\\ &\qquad\qquad\cdot\Big{(}\nu_{k}\wedge\nu_{\ell}+\nu_{\ell}\wedge\bar{\nu}_{k}+\bar{\nu}_{k}\wedge\nu_{\ell}\,|\mu_{F}|^{4}+\bar{\nu}_{\ell}\wedge\nu_{k}\,|\mu_{F}|^{4}\Big{)}\,d^{2}z\\ =&-2\sqrt{-1}\int_{\mathbb{D}}(1+|\mu_{F}|^{2})\,e^{\phi}\,\Big{(}\nu_{k}\wedge\bar{\nu}_{\ell}+\nu_{\ell}\wedge\bar{\nu}_{k}\Big{)}\,d^{2}z.\end{split}$
@@
(5.14) $-\sqrt{-1}\,i^{*}_{+}\omega_{C}=-2\int_{\mathbb{D}}(1+|\mu_{F}|^{2})\,e^{\phi}\,\sum_{k,\ell\in\mathbb{Z}}\Big{(}\nu_{k}\wedge\bar{\nu}_{\ell}\Big{)}\,d^{2}z.$
@@
¯$\partial\bar{\partial}\Big{(}2\int_{\mathbb{D}}e^{\phi}|\mu_{F}|^{2}\,d^{2}z\Big{)}=-\sqrt{-1}\,i^{*}_{+}\omega_{\mathrm{C}}\qquad\text{over}\quad T_{0}(1)^{+}.$
@@
By Theorems 4.10 and 5.7, we have the following result:
@@
¯$\partial\bar{\partial}\,\big{(}2E\big{)}=\partial\bar{\partial}\Big{(}2\int_{\mathbb{D}}e^{\phi}|\mu_{F}|^{2}\,d^{2}z\Big{)}=\sqrt{-1}\,i^{*}_{\pm}\mathrm{Mess}^{*}_{\pm}\big{(}\omega_{\mathrm{WP}}\big{)}.$
@@
By Remark 4.13 and the proof of Theorem 5.7, the pullback 2-form $\mathrm{Mess}^{*}_{\pm}(\omega_{\mathrm{WP}})$ to $T_{0}(1)^{\pm}$ has the following expression:
(5.15) $i^{*}_{\pm}\mathrm{Mess}^{*}_{\pm}(\omega_{\mathrm{WP}})=2{\sqrt{-1}}\int_{\mathbb{D}}(1+|\mu_{F}|^{2})e^{\phi}\,\sum_{k,\ell\in\mathbb{Z}}\,\nu_{k}\wedge\bar{\nu}_{\ell}\ d^{2}z$
@@
Remark 5.10.
From equation (5.11), it follows that the holomorphic derivative of the anti-holomorphic energy functional $E$, associated with the harmonic map $F_{\pm}$, on $T_{0}(1)^{\pm}$ is given by the Hopf differential $\Phi(F_{\pm})$. In contrast, the holomorphic derivative of the universal Liouville action $S$ on $T_{0}(1)$ is expressed in terms of the Schwarzian of a univalent function on $\mathbb{D}$, determined by the conformal welding data, as established in Theorem 3.1 of [18]. Understanding the difference of these two holomorphic quadratic differentials is therefore essential for the analyzing the difference between $E$ and $S=\pi I_{L}$ , an issue that will be addressed in future work.